OpenStudy (anonymous):

find the equation for the line spanned by the vector u=(3,-2,5)

OpenStudy (nowhereman):

Any linear combination with this vector.

OpenStudy (amistre64):

we can use the vector itself a point; and calculate another point that the magnitude touches..

OpenStudy (amistre64):

or use 0,0,0 lol

OpenStudy (amistre64):

x=3t y=-2t z=5t

OpenStudy (amistre64):

or should we use a generic point (a,b,c)?

OpenStudy (amistre64):

(x,y,z) is prolly more standard x = x+3t; y=y-2t; z=z+5t